7 1018 MI what price do farmers get for their water elon crop s? In the third week of July, a random sample of 44 farming regions gave a sample mean of x·S6.88 per 100 pounds of watermelon. Assume that o is known to be $1.92 per 100 pounds. (a Find a 90% confidence interval for the population mean price per i00 pounds) that farmers in this region get for their watermelon crop what is the margin of error? (Round your answers to two decimal places.) lower limit $640 upper limit 736 margin of error $ 0.48 (b) Find the sample size necessary for a 90% confidence level with maximal error of estimate E-0.41 for the mean price per 100 pounds of watermelon (Round up to the nearest whole number.) farming regions to market. Find a 90% confidence interval for the population mean cash value of this crop, what is the margin of (c) A farm brings 15 tons of water error? Hint: 1 ton is 2000 pounds. (Round your answers to two decimal places.) lower limit upper limit $ 775 $6.01
Answer #1

(a) here sample mean \bar x = 6.88 pounds per 100

standard deviation =\sigma= $ 1.92 pounds per 100

sample size = n = 44

standard error of sample means = se0 = \sigma /sqrt(n) = 1.92/sqrt(44) = 0.2895

90% confidence interval = \bar x +- Zcritical  se0 = 6.88 +- 1.645 * 0.2895

= (6.404 , 7.356)

(b) Here maximum margin of estimate E = 0.41

so margin of error = critical test statistic * standard error of sample mean

0.41 = 1.645 * 1.92/sqrt(n)

sqrt(n) = 59.34 or 60

(c) Here 1 ton = 2000 pounds

total crop = 15 tonnes =30000 pounds

so here expected mean cash value of this croup = 300 * $ 6.88 = $ 2064

standad deviation of population mean cash value = $ 1.92 * 300 = $ 576

Lower limit = 2064 - 1.645 * 576/sqrt(300) = $ 2009.29

Upper limit = 2064 + 1.645 * 576/sqrt(300) = $ 2118.71

margin of error = $ 109.92

Here 90% confidence interval =